Hi:u=85 Ha:u=85




The sample mean is 83,the sample size is 39,and the standard deviation is 5. Use a=0.05

1
Expert's answer
2022-05-26T11:06:52-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=85"

"H_1:\mu\not=85"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=38" and the critical value for a two-tailed test is "t_c =2.024394."

The rejection region for this two-tailed test is "R = \{t:|t|>2.024394\}."

The t-statistic is computed as follows:



"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{83-85}{5/\sqrt{39}}=-2.498"

Since it is observed that "|t|=2.498>2.024394=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed, "df=38" degrees of freedom, "t=-2.498" is "p= 0.016935," and since "p= 0.016935<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 85, at the "\alpha = 0.05" significance level.